# Deconfinement and chiral phase transitions in quark matter with chiral imbalance

### v1

@@ Abstract
We study the thermodynamics of the Polyakov–Nambu–Jona-Lasinio model considering the effects of an effective chiral chemical potential. We offer a new parametrization of the Polyakov loop potential depending on temperature and the chemical potential which, when used together with a proper regularization scheme of vacuum contributions, predicts results consistent with those from lattice simulations.
@@ I Introduction
The effects of a chiral imbalance in a medium can be implemented in the grand canonical ensemble by introducing a chiral chemical potential $\mu_{5}$. Chiral asymmetry effects on quark matter and applications for the quantum chromodynamics (QCD) phase diagram dualities were developed in Refs. Khunjua:2021oxf ; Khunjua:2018jmn ; Khunjua:2020vrp ; Khunjua:2018sro . The influence of external electric and magnetic fields and $\mu_{5}$ on the QCD chiral phase transition was investigated in Ruggieri:2016xww . As far applications to QCD are concerned, one very interesting aspect is related to the fact that QCD at finite chiral chemical potential is amenable to lattice simulations, since it is free from the sign problem, contrary to what happens, for example, for the case of a finite baryon chemical potential. In the lattice studies that implemented the effect of a chiral chemical potential Braguta:2015owi ; Braguta:2015zta it was predicted in particular that the critical temperature ($T_{c}$) for chiral symmetry restoration increases with $\mu_{5}$. However, the predictions from the literature based on effective models like the Nambu–Jona-Lasinio (NJL) model Ruggieri:2011xc ; Fukushima:2010fe ; Chao:2013qpa ; Yu:2014sla ; Yu:2015hym and the quark linear sigma models Ruggieri:2011xc ; Chernodub:2011fr have found exactly the opposite behavior, i.e., those model studies have shown a critical temperature which is a decreasing function of $\mu_{5}$ instead. It was argued in Ref. Braguta:2016aov that $\mu_{5}$ can favor quark-antiquark pairing, increasing the quark condensate and, consequently, requiring a higher temperature for the chiral symmetry restoration, which is consistent with the results predicted in Refs. Braguta:2015owi ; Braguta:2015zta . An agreement between the behavior for $T_{c}$ as a function of $\mu_{5}$ as predicted by lattice simulations was obtained by universality arguments in the large $N_{c}$ limit (where $N_{c}$ is the number of color degrees of freedom) in Ref. Hanada:2011jb , in studies with phenomenological quark-gluon interactions in the framework of the Dyson-Schwinger equations Wang:2015tia ; Xu:2015vna ; Shi:2020uyb ; Tian:2015rla ; Cui:2016zqp , nonlocal NJL models Frasca:2016rsi ; Ruggieri:2016ejz ; Ruggieri:2020qtq , through the use of a self-consistent mean field approximation in the NJL model Yang:2020ykt ; Liu:2020elq ; Liu:2020fan ; Yang:2019lyn ; Pan:2016ecs ; Lu:2016uwy , using a nonstandard renormalization scheme in the quark linear sigma model Ruggieri:2016cbq , and also in the context of chiral perturbation theory Espriu:2020dge ; GomezNicola:2023ghi ; Andrianov:2019fwz .
@@
In this work, we extend the methodology introduced in Ref. Farias:2016let to explore the dynamics of the confinement-deconfinement transition influenced by $\mu_{5}$ within the Polyakov extended NJL model (PNJL). This investigation aligns with predictions from lattice QCD Braguta:2015owi , where the critical temperatures associated with the confinement-deconfinement transition and the chiral symmetry restoration increase with $\mu_{5}$. Notably, to the best of our knowledge, there is no framework that simultaneously captures the behavior of both (pseudo) criticial temperatures in accordance with lattice predictions. As we show in this work, the agreement of the results requires the application of the MSS procedure once again but also a new parametrization of the Polyakov loop potential in order to include the effect of the chiral chemical potential besides of the standard one given in terms of the temperature only. In this work, we offer such an appropriate parametrization to reproduce the lattice results.
The remainder of this paper is organized as follows. In Sec. II we discuss the PNJL model studied in this work along also with the applied regularization scheme. In Sec. III we present our numerical results. We compare our regularization scheme with the one more commonly considered in the literature. We propose a new parameterization of the Polyakov loop, depending on $T$ and $\mu_{5}$, which turns out to be more appropriate to describe the effects due to a chiral chemical potential. A detailed analysis of the phase diagram of the model is presented in Sec. IV. Finally, our conclusions and final remarks are presented in Sec. V.
@@ II.1 Regularization
In this work, we compare the results obtained by the traditional three-dimensional cutoff regularization scheme (TRS), and the medium separation scheme (MSS) that have been frequently used in the literature to describe the QCD phase structure in different contexts Carlomagno:2021gcy ; Schaefer:2009ui ; Carlomagno:2021gcy ; Braguta:2015owi ; Lopes:2021tro . The MSS regularization was first proposed in Ref. Farias:2005cr , and was first applied to a problem involving a chiral imbalance in Ref. Farias:2016let . In the following, we summarize the implementation of this regularization scheme. We start from the mass gap equation at $T=\mu=0$:
@@
$\frac{1}{x^{2}+\omega_{s}^{2}}$ $=\frac{1}{x^{2}+p^{2}+M^{2}}$
$+\frac{p^{2}+M_{0}^{2}-\omega_{s}^{2}}{(x^{2}+p^{2}+M^{2})(x^{2}+\omega_{s}^{2})},$ (13)
@@
$-\frac{2\mu_{5}^{2}+M^{2}-M_{0}^{2}}{8\pi^{2}}+\frac{M^{2}-2\mu_{5}^{2}}{8\pi^{2}}\left(\frac{M^{2}}{M_{0}^{2}}\right).$ (23)
@@
To motivate and emphasize the importance of the correct separation of medium contributions in the presence of $\mu_{5}$, we compare the results predicted by both schemes for the chiral density $\langle n_{5}\rangle$. From chiral perturbation theory (ChPT), the QCD partition function in the presence of $\mu_{5}$ is modified to Astrakhantsev:2019wnp :
@@
which is a linear function of $\mu_{5}$ and agrees with the linear behavior predicted by ChPT. Moreover, the slope is precisely the one predicted by ChPT, $4f_{\pi}^{2}$, as we show next. To this end, we follow the works related to the origin of the MSS, referred to as Implicit Regularization Scheme Battistel:1998tj ; Farias:2005cr . Specifically, the authors of those references showed that
@@
at an arbitrary mass scale $M$, which can still be a function of $\mu_{5}$ [^1] In Ref. Farias:2005cr the authors were interested in the color superconducting gap $\Delta$ behavior, but their approach may be used for the mass scale $M$ without loss of generality.. We note that $4i\tilde{I}_{\text{log}}(x^{2})=I_{\text{log}}(x^{2})$. This definition is valid at any mass scale $M$, and can be expressed in terms of the vacuum quark mass, $M_{0}$, using the identity:
@@
This allows us to rewrite Eq. (29) as:
@@
With the use of $f_{\pi}^{2}$ from Eq. (30) and taking $N_{c}=3$, leads to:
@@ III Numerical Results at zero quark chemical potential
Recall that the correct order parameter for the chiral transition is the quark condensate, $\langle\bar{q}q\rangle$. However, given its linear relation with the effective quark mass, one may analyze the chiral symmetry restoration through the evolution of $M$ with the temperature and/or with the chemical potential without loss of generality. The phase diagrams in the $T\times\mu_{5}$ plane for the chiral and deconfinement transitions are shown in Fig. 4. The values of the pseudocritical temperatures at $\mu_{5}=0$, $T_{\rm pc}(0)$, are given in Tab. 2 for both regularization schemes.
@@
Recent lattice simulations predict a continuous and smooth increasing behavior of the order parameters with $\mu_{5}$ for both the chiral and deconfinement transitions. In addition, the values for $T_{\rm pc}^{c}$ and $T_{\rm pc}^{d}$ are approximately the same Braguta:2015owi . These results are represented by the squared dots in Fig. 4. However, one can see that the PNJL model in the presence of a chiral imbalance using the standard form for the fit of Polyakov loop, given by Eq. (5), is not able to reproduce the lattice QCD behavior. Although the MSS does not predict a critical end point (CEP) as the standard NJL model does Farias:2016let , both schemes lead to decreasing functions of the pseudocritical temperatures with $\mu_{5}$. We have also observed a similar behavior with other Polyakov loop parametrizations commonly considered in the literature. In all available parameterizations, even when using the MSS the pseudocritical temperatures are always decreasing functions of the chiral chemical potential. Our proposal here is that, in order to obtain results in agreement with the lattice results for the pseudocritical temperatures, these parameterizations need to be changed so that they not only depend on the temperature, but also depend on $\mu_{5}$. Thus, we propose an explicit redefinition of the Polyakov loop potential such as to include an explicit dependence of this function with $\mu_{5}$, i.e., $\mathcal{U}(\Phi,T)\to\mathcal{U}(\Phi,T,\mu_{5})$ [^2] As we have already mentioned previously, $\Phi=\Phi^{\dagger}$ at $\mu=0$.. Still working with the parametrization of the form in Eq. (7), we propose modifying the coefficient of $\Phi\Phi^{\dagger}$ in the potential $\mathcal{U}$ as follows:
@@
To obtain the MSS results presented in the remainder of this paper we have used $k_{1}=0.08602,k_{2}=-1.39646$ and $k_{3}=0.3381$. These values of constants can be seen to be the fitting values needed to reproduce the lattice results when a chiral chemical potential is present. Note that we have chosen a polynomial dependence on $\mu_{5}$ similar to the one on $T$ of the coeffecient $b_{2}$. We note that a plethora of different parametrization could be used in place of the polynomial we used. We have chosen the simplest one with the aim of showing that the missing ingredient for the models to reproduce lattice QCD simulations is a dependence of the Polyakov loop potential with $\mu_{5}$, in addition to the dependence on $T$.
Figure 5: MSS results for the normalized quark mass $M/M_{0}$ on panel (a) and Polyakov loop $\Phi$ on panel (b) as functions of the temperature, for $\mu=0$ and different values of $\mu_{5}$. Results obtained with the redefined Polyakov loop potential as given by Eq. (36).
@@
Figure 6: Phase diagrams in the plane $T\times\mu_{5}$ for chiral restoration and deconfinement, comparing lattice results from Ref. Braguta:2015owi and MSS, including a dependence with $\mu_{5}$ on the Polyakov loop potential, as given in Eq. (36). The dotted and dashed lines are the MSS results for chiral and deconfinement transitions, respectively, while large dots are lattice results.
@@ III.1 Thermodynamic quantities
It is common to find in the literature an alternative definition of the squared speed of sound, calculated at fixed entropy per baryon $s/n_{B}$. This alternative definition is motivated in the context of heavy ion collisions since the speed of sound is approximately constant in the whole expansion stage of the collision Braun-Munzinger:2015hba . Since we are working at zero quark density, we will use the definition of $c_{s}^{2}$ at finite chiral chemical potential as given in Eq. (42). Finally, we also consider the scaled trace anomaly $\Delta$, also referred to in the literature as “interaction measure”:
@@
Results for $p_{N}$, $s$, $\epsilon$, $c^{2}_{s}$, and $\Delta$ are shown in Figs 7 and 8. Initially, it is worth mentioning that the results for these thermodynamic quantities converge to the Stefan-Boltzmann limit, i.e. the limit of a gas of noninteracting particles. We recall that the Stefan-Boltzmamm limit of the pressure is given by Costa:2009ae :
@@
It is known that the scaled trace anomaly $\Delta$ has peaks sharply at the deconfined temperature and has a tail that approaches zero at asymptotically large values of $T$ in pure gauge theories. A similar behavior is obtained for the PNJL model as shown in Fig. 8 for both methods, with the caveat that at $\mu_{5}=0.5$ GeV the TRS curve diverges at the critical temperature. Also in Fig. 8, we show a comparison of the TRS and MSS results for the equation of state (EoS) and the squared speed of sound. From Eq. (42), one can see that $c_{s}^{2}$ represents the slope of the curve $p_{n}\times\varepsilon$ at each temperature value. Once again, the main difference between the curves occurs at $\mu_{5}=0.5$ GeV for TRS, for which the speed of sound goes to zero in the first-order transition region. To express the pressure and energy density for the same temperature range used for $c_{s}^{2}$ it is more convenient to use a logarithmic scale in the EoS plot. This is the reason why the jump at $p_{N}\lesssim$ 0.1 GeV/fm3 appears to be more pronounced than the equivalent discontinuity in the squared speed of sound. On the other hand, the MSS curve does not show a first-order transition; the step-like behavior in the region $0.2\leavevmode\nobreak\ {\rm GeV}\lesssim T\lesssim 0.3\leavevmode\nobreak\ {\rm GeV}$ is related to a continuous change of slope of the EoS in the range $0.8\leavevmode\nobreak\ {\rm GeV/fm}^{3}\lesssim\varepsilon\lesssim 10\leavevmode\nobreak\ {\rm GeV/fm}^{3}$. For the other values of $\mu_{5}$ in both methods, the curves approach the conformal limit of 1/3 at values of $T$ of the order of $0.6$ GeV.

### v2

@@ Abstract
We study the thermodynamics of the Polyakov-Nambu-Jona-Lasinio model considering the effects of an effective chiral chemical potential. We offer a new parametrization of the Polyakov-loop potential depending on temperature and the chiral chemical potential which, when used together with a proper regularization scheme of vacuum contributions, predicts results consistent with those from lattice simulations.
@@ I Introduction
The effects of a chiral imbalance in a medium can be implemented in the grand canonical ensemble by introducing a chiral chemical potential $\mu_{5}$. Chiral asymmetry effects on quark matter and applications for the quantum chromodynamics (QCD) phase diagram dualities were developed in Refs. Khunjua:2021oxf ; Khunjua:2018jmn ; Khunjua:2020vrp ; Khunjua:2018sro . The influence of external electric and magnetic fields and $\mu_{5}$ on the QCD chiral phase transition was investigated in Ruggieri:2016xww . As far applications to QCD are concerned, one very interesting aspect is related to the fact that QCD at finite chiral chemical potential is amenable to lattice simulations, since it is free from the sign problem, contrary to what happens, for example, for the case of a finite baryon chemical potential. In the lattice studies that implemented the effect of a chiral chemical potential Braguta:2015owi ; Braguta:2015zta it was predicted in particular that the critical temperature ($T_{c}$) for chiral symmetry restoration increases with $\mu_{5}$. However, the predictions from the literature based on effective models like the Nambu-Jona-Lasinio (NJL) model Ruggieri:2011xc ; Fukushima:2010fe ; Chao:2013qpa ; Yu:2014sla ; Yu:2015hym and the quark linear sigma models Ruggieri:2011xc ; Chernodub:2011fr have found exactly the opposite behavior, i.e., those model studies have shown a critical temperature which is a decreasing function of $\mu_{5}$ instead. It was argued in Ref. Braguta:2016aov that $\mu_{5}$ can favor quark-antiquark pairing, increasing the quark condensate and, consequently, requiring a higher temperature for the chiral symmetry restoration, which is consistent with the results predicted in Refs. Braguta:2015owi ; Braguta:2015zta . An agreement between the behavior for $T_{c}$ as a function of $\mu_{5}$ as predicted by lattice simulations was obtained by universality arguments in the large $N_{c}$ limit (where $N_{c}$ is the number of color degrees of freedom) in Ref. Hanada:2011jb , in studies with phenomenological quark-gluon interactions in the framework of the Dyson-Schwinger equations Wang:2015tia ; Xu:2015vna ; Shi:2020uyb ; Tian:2015rla ; Cui:2016zqp , nonlocal NJL models Frasca:2016rsi ; Ruggieri:2016ejz ; Ruggieri:2020qtq , through the use of a self-consistent mean field approximation in the NJL model Yang:2020ykt ; Liu:2020elq ; Liu:2020fan ; Yang:2019lyn ; Pan:2016ecs ; Lu:2016uwy , using a nonstandard renormalization scheme in the quark linear sigma model Ruggieri:2016cbq , and also in the context of chiral perturbation theory Espriu:2020dge ; GomezNicola:2023ghi ; Andrianov:2019fwz .
@@
In this work, we extend the methodology introduced in Ref. Farias:2016let to explore the dynamics of the confinement-deconfinement transition influenced by $\mu_{5}$ within the Polyakov extended NJL model (PNJL). This investigation aligns with predictions from lattice QCD Braguta:2015owi , where the critical temperatures associated with the confinement-deconfinement transition and the chiral symmetry restoration increase with $\mu_{5}$. Notably, to the best of our knowledge, there is no framework that simultaneously captures the behavior of both (pseudo) criticial temperatures in accordance with lattice predictions. As we show in this work, the agreement of the results requires the application of the MSS procedure once again but also a new parametrization of the Polyakov-loop potential in order to include the effect of the chiral chemical potential besides of the standard one given in terms of the temperature only. In this work, we offer such an appropriate parametrization to reproduce the lattice results.
The remainder of this paper is organized as follows. In Sec. II we discuss the PNJL model studied in this work along also with the applied regularization scheme. In Sec. III we present our numerical results. We compare our regularization scheme with the one more commonly considered in the literature. We propose a new parametrization of the Polyakov loop, depending on $T$ and $\mu_{5}$, which turns out to be more appropriate to describe the effects due to a chiral chemical potential. A detailed analysis of the phase diagram of the model is presented in Sec. IV. Finally, our conclusions and final remarks are presented in Sec. V.
@@ II.1 Regularization
In this work, we compare the results obtained by the traditional three-dimensional cutoff regularization scheme (TRS), and the MSS that have been frequently used in the literature to describe the QCD phase structure in different contexts Carlomagno:2021gcy ; Schaefer:2009ui ; Carlomagno:2021gcy ; Braguta:2015owi ; Lopes:2021tro . The MSS regularization was first proposed in Ref. Farias:2005cr , and was first applied to a problem involving a chiral imbalance in Ref. Farias:2016let . In the following, we summarize the implementation of this regularization scheme. We start from the mass gap equation at $T=\mu=0$:
@@
$\frac{1}{x^{2}+\omega_{s}^{2}}$ $=\frac{1}{x^{2}+p^{2}+M_{0}^{2}}$
$+\frac{p^{2}+M_{0}^{2}-\omega_{s}^{2}}{(x^{2}+p^{2}+M_{0}^{2})(x^{2}+\omega_{s}^{2})},$ (13)
@@
$-\frac{2\mu_{5}^{2}+M^{2}-M_{0}^{2}}{8\pi^{2}}+\frac{M^{2}-2\mu_{5}^{2}}{8\pi^{2}}\ln{\left(\frac{M^{2}}{M_{0}^{2}}\right)}.$ (23)
@@
To motivate and emphasize the importance of the correct separation of medium contributions in the presence of $\mu_{5}$, we compare the results predicted by both schemes for the chiral density $\langle n_{5}\rangle$. From chiral perturbation theory (ChPT), the QCD partition function in the presence of $\mu_{5}$ is modified to Astrakhantsev:2019wnp
@@
which is a linear function of $\mu_{5}$ and agrees with the linear behavior predicted by ChPT. Moreover, the slope is precisely the one predicted by ChPT, $4f_{\pi}^{2}$, as we show next. To this end, we follow the works related to the origin of the MSS, referred to as the Implicit Regularization Scheme Battistel:1998tj ; Farias:2005cr . Specifically, the authors of those references showed that
@@
at an arbitrary mass scale $M$, which can still be a function of $\mu_{5}$. [^1] In Ref. Farias:2005cr the authors were interested in the color superconducting gap $\Delta$ behavior, but their approach may be used for the mass scale $M$ without loss of generality. We note that $4i\tilde{I}_{\text{log}}(x^{2})=I_{\text{log}}(x^{2})$. This definition is valid at any mass scale $M$, and can be expressed in terms of the vacuum quark mass, $M_{0}$, using the identity,
@@
This allows us to rewrite Eq. (29) as
@@
With the use of $f_{\pi}^{2}$ from Eq. (30) and taking $N_{c}=3$, leads to
@@ III Numerical Results at zero quark chemical potential
Recall that the correct order parameter for the chiral transition is the quark condensate, $\langle\bar{q}q\rangle$. However, given its linear relation with the effective quark mass, one may analyze the chiral symmetry restoration through the evolution of $M$ with the temperature and/or with the chemical potential without loss of generality. The phase diagrams in the $T\times\mu_{5}$ plane for the chiral and deconfinement transitions are shown in Fig. 4. The values of the pseudocritical temperatures at $\mu_{5}=0$, $T_{\rm pc}(0)$, are given in Table 2 for both regularization schemes.
@@
Recent lattice simulations predict a continuous and smooth increasing behavior of the order parameters with $\mu_{5}$ for both the chiral and deconfinement transitions. In addition, the values for $T_{\rm pc}^{c}$ and $T_{\rm pc}^{d}$ are approximately the same Braguta:2015owi . These results are represented by the squared dots in Fig. 4. However, one can see that the PNJL model in the presence of a chiral imbalance using the standard form for the fit of Polyakov loop, given by Eq. (5), is not able to reproduce the lattice QCD behavior. Although the MSS does not predict a critical end point (CEP) as the standard NJL model does Farias:2016let , both schemes lead to decreasing functions of the pseudocritical temperatures with $\mu_{5}$. We have also observed a similar behavior with other Polyakov-loop parametrizations commonly considered in the literature. In all available parametrizations, even when using the MSS the pseudocritical temperatures are always decreasing functions of the chiral chemical potential. Our proposal here is that, in order to obtain results in agreement with the lattice results for the pseudocritical temperatures, these parametrizations need to be changed so that they not only depend on the temperature, but also depend on $\mu_{5}$. Thus, we propose an explicit redefinition of the Polyakov-loop potential such as to include an explicit dependence of this function with $\mu_{5}$, i.e., $\mathcal{U}(\Phi,T)\to\mathcal{U}(\Phi,T,\mu_{5})$. [^2] As we have already mentioned previously, $\Phi=\Phi^{\dagger}$ at $\mu=0$. Still working with the parametrization of the form in Eq. (7), we propose modifying the coefficient of $\Phi\Phi^{\dagger}$ in the potential $\mathcal{U}$ as follows:
@@
To obtain the MSS results presented in the remainder of this paper we have used $k_{1}=0.08602,k_{2}=-1.39646$ and $k_{3}=0.3381$. These values of constants can be seen to be the fitting values needed to reproduce the lattice results when a chiral chemical potential is present.
Traditional implementations of the Polyakov loop in quark models typically rely on fitting lattice results for a gluon gas at finite temperature. However, these models often fail to accurately replicate the qualitative behavior observed in lattice QCD simulations, indicating that both the quark and gauge sectors are influenced by medium effects beyond temperature. To address this discrepancy, it is common to introduce additional dependencies into the Polyakov-loop potential to align with lattice data. For example, the authors in Ref. Schaefer:2007pw proposed a dependence on quark chemical potential $\mu$ and the number of flavors, employing renormalization group arguments to investigate the thermodynamic properties of the quark-meson model. Similarly, Ref. Adhikari:2018cea modified the dependence on $\mu$ by introducing the isospin chemical potential $\mu_{I}$. In Ref. Ferreira:2013tba it was incorporated an implicit dependence on the magnetic field in the entangled PNJL (EPNJL) model, where the coupling constant is modified to depend on the order parameters $\Phi$ and $\bar{\Phi}$. The resulting model then predicted inverse magnetic catalysis, a feature observed in lattice simulations but not in the ordinary PNJL approach. The authors of Ref. Sun:2024anu studied rotating systems by including a dependence on angular velocity $\omega$ in $\cal U$ and $\Phi$, finding that critical temperatures increase with $\omega$ for both chiral and deconfinement transitions, consistent with lattice data. Then, in Ref. Dexheimer:2009hi it was proposed a dependence of the chemical potential on the Polyakov-loop potential that persists at $T=0$, enabling the study of density-driven deconfinement. This possibility is particularly relevant in light of recent gravitational wave data analyses.
In this study, we adopted a polynomial dependence of the coefficient $\bar{b}_{2}$ on $\mu_{5}$, analogous to its existing dependence on $T$ as shown in Eq. (37). We acknowledge that numerous alternative parametrizations could be employed instead of the chosen polynomial. However, we opted for the simplest approach to demonstrate that the key missing element for these models to accurately replicate lattice QCD simulations is a dependence of the Polyakov-loop potential on $\mu_{5}$, in addition to its dependence on $T$.
Figure 5: MSS results for the normalized quark mass $M/M_{0}$ on panel (a) and Polyakov loop $\Phi$ on panel (b) as functions of the temperature, for $\mu=0$ and different values of $\mu_{5}$. Results obtained with the redefined Polyakov-loop potential as given by Eq. (36).
@@
Figure 6: Phase diagrams in the plane $T\times\mu_{5}$ for chiral restoration and deconfinement, comparing lattice results from Ref. Braguta:2015owi and MSS, including a dependence with $\mu_{5}$ on the Polyakov-loop potential, as given in Eq. (36). The dotted and dashed lines are the MSS results for chiral and deconfinement transitions, respectively, while large dots are lattice results.
@@ III.1 Thermodynamic quantities
It is common to find in the literature an alternative definition of the squared speed of sound, calculated at fixed entropy per baryon $s/n_{B}$. This alternative definition is motivated in the context of heavy ion collisions since the speed of sound is approximately constant in the whole expansion stage of the collision Braun-Munzinger:2015hba . Since we are working at zero quark density, we will use the definition of $c_{s}^{2}$ at finite chiral chemical potential as given in Eq. (42). Finally, we also consider the scaled trace anomaly $\Delta$, also referred to in the literature as ‘interaction measure’:
@@
Results for $p_{N}$, $s$, $\epsilon$, $c^{2}_{s}$, and $\Delta$ are shown in Figs 7 and 8. Initially, it is worth mentioning that the results for these thermodynamic quantities converge to the Stefan-Boltzmann limit, i.e. the limit of a gas of noninteracting particles. We recall that the Stefan-Boltzmamm limit of the pressure is given by Costa:2009ae
@@
It is known that the scaled trace anomaly $\Delta$ has peaks sharply at the deconfined temperature and has a tail that approaches zero at asymptotically large values of $T$ in pure gauge theories. A similar behavior is obtained for the PNJL model as shown in Fig. 8 for both methods, with the caveat that at $\mu_{5}=0.5$ GeV the TRS curve diverges at the critical temperature. Also in Fig. 8, we show a comparison of the TRS and MSS results for the equation of state (EOS) and the squared speed of sound. From Eq. (42), one can see that $c_{s}^{2}$ represents the slope of the curve $p_{n}\times\varepsilon$ at each temperature value. Once again, the main difference between the curves occurs at $\mu_{5}=0.5$ GeV for TRS, for which the speed of sound goes to zero in the first-order transition region. To express the pressure and energy density for the same temperature range used for $c_{s}^{2}$ it is more convenient to use a logarithmic scale in the EOS plot. This is the reason why the jump at $p_{N}\lesssim$ 0.1 GeV/fm3 appears to be more pronounced than the equivalent discontinuity in the squared speed of sound. On the other hand, the MSS curve does not show a first-order transition; the steplike behavior in the region $0.2\leavevmode\nobreak\ {\rm GeV}\lesssim T\lesssim 0.3\leavevmode\nobreak\ {\rm GeV}$ is related to a continuous change of slope of the EOS in the range $0.8\leavevmode\nobreak\ {\rm GeV/fm}^{3}\lesssim\varepsilon\lesssim 10\leavevmode\nobreak\ {\rm GeV/fm}^{3}$. For the other values of $\mu_{5}$ in both methods, the curves approach the conformal limit of 1/3 at values of $T$ of the order of $0.6$ GeV.